If you travel 700ft every ten seconds. How many miles per hour are you traveling?
Answer by nyc_function(2741)

There are 3600 in one hour.
So, 10/3600 = (1/360) of an hour.
D = rt
D = distance
r = rate in mph
t = time (in hours)
700 = (1/360)r
700 divided by (1/360) = r
252,000 = r
The miles per hour is 252,000.
We now muct convert 252,000 feet per hour to miles per hour.
252000ft per hour / 5280ft per mile = 47.72727272... miles per hour
Understand?